Hello.
I have a sequence : 4 8 28 108 608 3108
I can't solve it.
How to do it?

Answer by addingup(3677)

The beginning of your sequence follows the pattern of alternate 3^n+1 and 3^n-1. But then you go 108, 608, 3108 and that pattern, if there is one, I have not yet figured out.
:
3^1+1 = 4
3^2-1 = 8
3^3+1 = 28
If you follow this pattern, the next three numbers would be:
3^4-1 = 80
3^5+1 = 244
3^6-1 = 728